OpenStudy (anonymous):

givin that cot m=0.75 and cos m<0 what is the value of sin m

OpenStudy (amistre64):

cot = 3/4 tan = 4/3 3-4-5 triangle sin = 4/5 but is in Q3, sin = -4/5